You burn 8.0 g of marshmallow to warm 400 mL of water. The temperature of the water increases from 20 degrees C to 80 degrees C


How many calories of heat are transferred to the water?


What is the calories per gram of marshmallow?


The label on the bag of marshmallows indicates that one 8-gram marshmallow has 24 kcal. How does this compare with your answer. Explain.




1
Expert's answer
2021-05-27T07:34:14-0400

According to the heat equation:

Q = cm(T2 - T1)

where Q - released/absorbed heat, c - heat capacity, m - mass, T2 - final temperature, T1 - initial temperature.

Mass of water equals:

m = volume × density = 400 ml × 1 g/ml = 400 g = 0.4 kg

From here, the heat absorbed by water equals:

Q = 4182 J/kg°C × 0.4 kg × (80 °C - 20 °C) = 100368 J = 23988.5 cal = 24 kcal

As a result, the absorbed heat and the indicated energy are equal showing that all energy released from the burning marshmallow was directed to heat water.